# RetroGFN: Diverse and Feasible Retrosynthesis using GFlowNets

### v1

@@ Abstract
Single-step retrosynthesis aims to predict a set of reactions that lead to the creation of a target molecule, which is a crucial task in molecular discovery. Although a target molecule can often be synthesized with multiple different reactions, it is not clear how to verify the feasibility of a reaction, because the available datasets cover only a tiny fraction of the possible solutions. Consequently, the existing models are not encouraged to explore the space of possible reactions sufficiently. In this paper, we propose a novel single-step retrosynthesis model, RetroGFN, that can explore outside the limited dataset and return a diverse set of feasible reactions by leveraging a feasibility proxy model during the training. We show that RetroGFN achieves competitive results on standard top-k accuracy while outperforming existing methods on round-trip accuracy. Moreover, we provide empirical arguments in favor of using round-trip accuracy which expands the notion of feasibility with respect to the standard top-k accuracy metric.
@@ 1 Introduction
The rising interest in machine learning led to the development of many deep generative models for de novo drug design [1, 2, 3, 4, 5]. Such approaches can propose novel molecules with promising properties (e.g. high binding affinity score) predicted by other machine learning models [6, 7, 8, 9, 10], however, these virtual compounds eventually need to be synthesized and evaluated in the wet lab. This motivates the development of reliable (retro)synthesis planning algorithms able to design a synthesis route for an input molecule. Retrosynthesis aims to recursively decompose a target compound into simpler molecules forming a synthesis tree. The leaves of the tree are purchasable molecules from which the synthesis process can start and the tree itself is a synthesis recipe. By going bottom-up the tree and performing reactions defined by the tree nodes, one will eventually obtain the target molecule. The construction of such a tree usually consists of two components: a single-step retrosynthesis model that decomposes a molecule [11, 12, 13, 14, 15, 16], and a multi-step planning algorithm that guides the recursive decomposition to obtain the full synthesis tree [17, 18, 16, 19, 20, 21, 22, 23]. In this paper, we focus on single-step retrosynthesis, which predicts a reaction that is likely to synthesize a given molecule.
@@
In practice, many feasible reactions can lead to a given product. Since the success of a synthesis plan depends on factors that may vary over time (e.g. the availability or cost of reactants), the retrosynthesis model should ideally return all possible reactions. In other words, we would like to produce a diverse set of feasible reactions leading to the requested product. However, the available datasets cover only a fraction of feasible reactions, so for many of the included products, a lot of alternative reactions are missing. This limitation of current reaction datasets causes two major issues that we address in this paper.
First, the typical evaluation of retrosynthesis models involves the use of top-k accuracy, which verifies how many top-k reactions returned by the model are included in a given dataset. Our analysis performed on the USPTO-50k test split [24, 12, 25] reveals that on average more than 100 feasible reactions returned by the examined retrosynthesis models are ignored by top-k accuracy (see Figure 2). Since it is practically impossible to include all possible reactions in a finite dataset, one remedy relies on employing a machine learning model, which reliably assesses the reaction feasibility. This approach is applied in round-trip accuracy, a less exploited alternative to the top-k accuracy metric. Our experimental study demonstrates that replacing top-k accuracy with top-k round-trip accuracy decreases the number of ignored reactions to less than 20 elements (Figure 2) while being robust to non-trivially unfeasible reactions (Figure 2). In consequence, round-trip accuracy should be taken into account as a complementary metric to standard top-k accuracy. We recommend reporting it in future papers to make model evaluation more comprehensive (see Section 3 for detailed analysis).
@@
Second, since typical evaluation relies on using top-k accuracy (instead of round-trip accuracy), the existing retrosynthesis models are not encouraged to explore the space of feasible reactions well. Taking inspiration from the construction of round-trip accuracy, we employ a machine learning model, which rewards the retrosynthesis model for returning highly feasible reactions (not only those included in a fixed dataset). The main contribution of the paper is the development of a RetroGFN model that can explore beyond the dataset and return a diverse set of feasible reactions (Section 4). RetroGFN is based on the recent GFlowNet framework [26, 27] which enables exploration of the solution space and sampling from that space with probability proportional to the reward function, e.g., reaction feasibility. In consequence, GFlowNets can sample a large number of highly scored and diverse solutions. Our RetroGFN model leverages this property, sampling a large number of feasible reactions. It outperforms existing methods on the round-trip accuracy metric while achieving competitive results on the top-k accuracy.
@@
  • We provide empirical arguments for the importance of reporting the round-trip accuracy in the single-step retrosynthesis model evaluation (Section 3).
  • We develop RetroGFN: a model based on the GFlowNet framework that generates diverse and feasible reactions. To our knowledge, we are the first to adapt GFlowNets for retrosynthesis (Section 4). We make the code publicly available [^1] https://github.com/gmum/RetroGFN.
  • We benchmark the state-of-the-art single-step retrosynthesis models and show that our RetroGFN outperforms all considered models on the round-trip accuracy while achieving competitive results on the top-k accuracy (Section 5).
@@ 2 Related Work
Single-step Retrosynthesis. The single-step retrosynthesis problem is well-known in the drug-discovery community. The methods in this field can be roughly divided into template-based and template-free. The former are based on reaction templates (also called rules, see Figure 3), which describe the graph-level transformation of molecules that are encountered in the reactions [11, 12, 14, 28]. Templates provide a strong inductive bias as they form a fixed set of possible transformations that the retrosynthesis model can perform. Template-free approaches, on the other hand, do not rely on a template and aim to generate the transformation of the product (the change of the bonds and atoms between reactants and product) [29, 30, 31, 32, 33] or generate the product from scratch [34, 16, 35, 36]. Our RetroGFN is a template-based model, but it is not limited to a fixed set of templates as it composes them using pre-defined patterns. The template composition process was inspired by RetroComposer [37] but remains substantially different: we implement the generation process in the GFlowNets framework; we use more general patterns; we parametrize the second phase to be order-invariant; we guarantee the second phase ends with product and reactant patterns that can be mapped; and finally, we map the atoms using a machine learning model (while RetroComposer uses a heuristic).
@@
GFlowNets. GFlowNets [27] are a type of generative methods devoted to sampling from high-dimensional distributions. GFlowNets were originally proposed as an alternative to MCMC (offering the benefits of amortization) and reinforcement learning (displaying a mode-seeking behavior, that is the ability to discover multiple diverse modes), and later shown to be equivalent with special cases of other generative methods [38, 39]. The diversity, in particular, is a desired property in multiple scientific discovery tasks [40], including molecule [26, 41, 42, 43], biological sequence [26, 44], crystal [45], conformer [46] and DNA-encoded library [47] generation tasks. Importantly, from a scientific discovery standpoint, GFlowNets have also been used in the active learning context, both in the multi-objective [44] and multi-fidelity [48] settings.

### v2

# Diverse and Feasible Retrosynthesis using GFlowNets

@@ Abstract
Single-step retrosynthesis aims to predict a set of reactions that lead to the creation of a target molecule, which is a crucial task in molecular discovery. Although a target molecule can often be synthesized with multiple different reactions, it is not clear how to verify the feasibility of a reaction, because the available datasets cover only a tiny fraction of the possible solutions. Consequently, the existing models are not encouraged to explore the space of possible reactions sufficiently. In this paper, we propose a novel single-step retrosynthesis model, RetroGFN, that can explore outside the limited dataset and return a diverse set of feasible reactions by leveraging a feasibility proxy model during the training. We show that RetroGFN achieves competitive results on standard top-k accuracy while outperforming existing methods on round-trip accuracy. Moreover, we provide empirical arguments in favor of using round-trip accuracy, which expands the notion of feasibility with respect to the standard top-k accuracy metric.
keywords:
GFlowNets , Retrosynthesis , Drug Discovery
journal: Information Sciences
[…]
@@ 1 Introduction
The rising interest in machine learning has led to the development of many deep generative models for de novo drug design [gomez2018automatic, maziarz2021learning]. Such approaches can propose novel molecules with promising properties (e.g., high binding affinity score) predicted by other machine learning models [walters2020applications, gainski2023chienn, chithrananda2020chemberta] , however, these virtual compounds eventually need to be synthesized and evaluated in the wet lab. This motivates the development of reliable (retro)synthesis planning algorithms able to design a synthesis route for an input molecule. Retrosynthesis aims to recursively decompose a target compound into simpler molecules, forming a synthesis tree. The leaves of the tree are purchasable molecules from which the synthesis process can start, and the tree itself is a synthesis recipe. By going bottom-up the tree and performing reactions defined by the tree nodes, one will eventually obtain the target molecule. The construction of such a tree usually consists of two components: a single-step retrosynthesis model that decomposes a molecule and a multi-step planning algorithm that guides the recursive decomposition to obtain the full synthesis tree [segler2018planning, coley2019robotic, schwaller2020predicting]. In this paper, we focus on single-step retrosynthesis, which predicts a reaction that is likely to synthesize a given molecule.
@@
In practice, many feasible reactions can lead to a given product. Since the success of a synthesis plan depends on factors that may vary over time (e.g., the availability or cost of reactants), the retrosynthesis model should ideally return all possible reactions. In other words, we would like to produce a diverse set of feasible reactions leading to the requested product. However, the available datasets cover only a fraction of feasible reactions, so for many of the included products, a lot of alternative reactions are missing. This limitation of current reaction datasets causes two major issues that we address in this paper.
@@
First, the existing retrosynthesis models are not encouraged to explore the space of feasible reactions well. The main contribution of the paper is the development of a RetroGFN model that can explore beyond the dataset and return a diverse set of feasible reactions. RetroGFN is based on the recent GFlowNet framework [bengio2021flow, bengio2023gflownet], which enables exploration of the solution space and sampling from that space with probability proportional to the reward function which we define using an auxiliary reaction feasibility prediction model to guide our RetroGFN training outside the limited reaction dataset. In consequence, our model samples a large number of feasible reactions. It outperforms existing methods on the round-trip accuracy metric while achieving competitive results on the top-k accuracy.
Second, the typical evaluation of retrosynthesis models involves the use of top-k accuracy, which verifies how many top-k reactions returned by the model are included in a given dataset. Our analysis performed on the USPTO-50k test split [schneider2016s, coley2017retrosim, jin2017predicting] reveals that, on average, more than 100 feasible reactions returned by the examined retrosynthesis models are ignored by top-k accuracy (see Figure 7). Since it is practically impossible to include all possible reactions in a finite dataset, one remedy relies on employing a machine learning model, which reliably assesses the reaction feasibility. This approach is applied in round-trip accuracy [schwaller2020predicting], a less exploited alternative to the top-k accuracy metric. Round-trip accuracy was shown to correlate more with human judgment than standard top-k accuracy [jaume2023transformer] and recommended as a complementary metric in previous works [westerlund2024chemformers, chen2021deep, graph2edit, igashov2023retrobridge]. In this paper, we additionally strengthen the arguments in favor of reporting round-trip accuracy in the papers and demonstrate that replacing top-k accuracy with top-k round-trip accuracy decreases the number of ignored reactions while being robust to non-trivially unfeasible reactions.
@@
  • We develop RetroGFN: a model based on the GFlowNet framework that generates diverse and feasible reactions. To our knowledge, we are the first to adapt GFlowNets for retrosynthesis. RetroGFN is guided by an auxiliary feasibility model that allows exploration outside the limited reaction dataset (Section 3). We make the code publicly available [^1] https://github.com/gmum/RetroGFN.
  • We benchmark the state-of-the-art single-step retrosynthesis models and show that our RetroGFN outperforms all considered models for $k>3$ on the round-trip accuracy while achieving competitive results on the top-k accuracy (Section 4).
  • We provide empirical arguments for the importance of reporting the round-trip accuracy in the single-step retrosynthesis model evaluation, which motivates our RetroGFN (Section 5).
@@ 2 Related Work
Single-step Retrosynthesis. The single-step retrosynthesis problem is well-known in the drug-discovery community. The methods in this field can be roughly divided into template-based, template-free, and semi-template. Template-based methods use reaction templates (also called rules, see Figure 1), which explicitly describe the graph-level transformation of molecules that are encountered in the reactions [segler2017neural, coley2017retrosim, dai2019retrosynthesis, baylon2019enhancing]. Templates provide a strong inductive bias as they form a fixed set of possible transformations that the retrosynthesis model can perform. Template-free approaches, on the other hand, do not rely on a template and aim to generate the transformation of the product (the change of the bonds and atoms between reactants and product) [sacha2021molecule, yan2020retroxpert, graphretro, wang2021retroprime] or generate the product from scratch [irwin2022chemformer, schwaller2020predicting, zheng2019predicting, mao2021molecular]. The semi-template models typically rely on the rules extracted from the training dataset, but without explicitly building a fixed set of reaction templates [graphretro, graph2edit]. Our RetroGFN is a semi-template model, as it is not limited to a fixed set of templates, but instead composes them using pre-defined patterns. The template composition process was inspired by RetroComposer [yan2022retrocomposer] but remains substantially different: we implement the generation process in the GFlowNets framework; we use more general patterns; we parametrize the second phase to be order-invariant; we guarantee the second phase ends with product and reactant patterns that can be mapped; and finally, we map the atoms using a machine learning model (while RetroComposer uses a heuristic).
@@
GFlowNets. GFlowNets [bengio2023gflownet] are a type of generative methods devoted to sampling from high-dimensional distributions. GFlowNets were originally proposed as an alternative to MCMC (offering the benefits of amortization) and reinforcement learning (displaying a mode-seeking behavior, that is, the ability to discover multiple diverse modes), and later shown to be equivalent to special cases of other generative methods [malkin2022gflownets]. The diversity, in particular, is a desired property in multiple scientific discovery tasks [bengio2021flow, jain2023multi, koziarski2024rgfn].
3 RetroGFN
RetroGFN is a single-step retrosynthesis model, meaning it predicts a set of molecules that could react to a given target product (see Figure 1 a)). A product is represented as an annotated graph $G=(V,T,E)$, where nodes $V=\{v_{1},v_{2},...,v_{n}\}$ correspond to the molecule’s atoms along with associated atom symbols (types) $T$, and edges $E$ are bonds. Additionally, each node and edge has an associated vector of features that will be used when embedding a molecule.
RetroGFN’s training is guided by a reaction feasibility model, enabling exploration outside the limited reaction dataset. More details on the guidance can be found in Section 3.3.
3.1 Reaction Templates and Patterns
(left side of the regular expression) and a set of reactants’ patterns (right side). The atoms of the product pattern are mapped to the atoms of the reactants’ patterns. Reaction templates provide a strong inductive bias to the model while limiting it to a fixed set of possible transformations.
However, we extend the covered reaction space by introducing a template composition process inspired by RetroComposer [yan2022retrocomposer]. In this approach, we choose the reaction center where the template is going to be applied and compose a concrete template step by step using the building blocks, called patterns.
We extract the templates from the train split of USPTO-50k, following [chen2021deep]. Each template is then split into product and reactant patterns (see Figure 1 b)). We denote a set of all encountered product patterns $PPS$ and an analogous set of reactant patterns $RPS$. The patterns do not include any molecular regular expression (SMARTS) and can be represented similarly to molecules, as annotated graphs.
3.2 Generation Process
  1. 1.
    The first phase determines a reaction center: a product pattern matched with the product.
  2. 2.
    The second phase gathers the reactant patterns.
  3. 3.
In the end, the obtained template is applied to the given product and results in a final set of reactants. Figure 2 shows an example of the composition process, while a detailed description of each phase can be found further in the section.
Figure 2: Illustration of the template composition process in RetroGFN for an input product. In the first phase, a product pattern and its concrete match to the atoms of the product are chosen. In the second phase, reactant patterns are gathered until all mappable atoms of the product pattern (highlighted red) can be mapped to mappable atoms of the reactant pattern (highlighted blue). In the third phase, the mapping between mappable products and reactant patterns is created, and the obtained template is applied, resulting in the reactants.
The core component of a GFlowNet model is a forward policy $P_{F}(a|s)$ describing the probability of taking action $a$ in the state $s$. The generation process samples a sequence of states and actions $\tau=(s_{1},a_{1},...,s_{k},a_{k},t)$ called a trajectory, where $t$ is a terminal state. In RetroGFN, an initial state $s_{1}$ is an input product, the intermediate states $s_{i}$ correspond to the partially constructed template, and the terminal state $t$ stores a final template along with a result of its application to the product. We group the states into three phases, and the specific definition of $P_{F}(a|s)$ depends on the phase $i$:
Score functions for all the phases share a common Graph Neural Network (GNN) encoder, denoted as $\mathrm{gnn}_{1}$ that given a product $p=(V,E,T)$, embeds its nodes’ features: $\mathrm{gnn}_{1}(p)\in\mathbb{R}^{n\times d}$, where $n$ is the number of product nodes and $d$ is the embedding size. We overload the notation and let $\mathrm{gnn}_{1}(v_{j})$ denote the embedding of a product node $v_{j}\in V$. The GNN architecture we use is similar to the one from LocalRetro: a stack of MPNN layers with a single Transformer layer [vaswani2017attention] on top. Details can be found in the A.
First Phase. A state $s$ in the first phase is an input product $p$. The action space $A^{1}(s)$ contains all possible atom matchings of product patterns from $PPS$ to the product $p$. An action $a\in A^{1}(s)$ contains the matched product pattern $pp\in PPS$ and the matched atom indices $I=\{i_{1},...,i_{m}\}$. The value of $i_{j}$ is an index of the product atom matched with $j$-th product pattern atom. To compute the $\mathrm{score}_{1}(s,a)$, we aggregate the representation of matched products’ nodes and put them into a multi-layer perceptron $\mathrm{MLP}_{1}:\mathbb{R}^{d}\rightarrow\mathbb{R}$:
Second Phase. The second phase iteratively adds reactant patterns to the
composed template. At the beginning of the phase, the list of reactant patterns is empty. The second phase action $a$ is a reactant pattern $rp_{j}\in RPS$ that is going to be added to the template. The $\mathrm{score}_{2}(s,a)$ concatenates the information from the previous phase and the reactant patterns collected so far (denoted as $R$) and feeds it to $\mathrm{MLP}_{2}:\mathbb{R}^{3d}\rightarrow\mathbb{R}^{|RPS|}$ that predicts the score for all the reactant patterns in $RPS$:
Here we select the $j$th score returned by the $\mathrm{MLP}_{2}$ as it corresponds to $rp_{j}$ reactant pattern from the action. Index embedding $e=\mathrm{E}_{A}(a)$ is a function that looks up the index of the element $a$ in the set $A$ and assigns the index a learnable embedding $e\in\mathbb{R}^{d}$ (e.g. $\mathrm{E}_{PPS}$ assign a unique learnable embedding to every $pp\in PPS$). At the end of this phase, we want to be sure that every atom from the product pattern can be mapped to some atom of the reactant pattern. Originally, each pattern had some atom mapping in the template it comes from (see Figure 1). Although those explicit mappings are inadequate in the novel-composed template, we can leverage the knowledge that an atom was originally mapped. For every pattern, we construct a set of mappable atoms that consists of the pattern’s atoms that were mapped in the original template (see Figure 3). The composed template is allowed to map only the mappable atoms. We ensure that all mappable atoms in the composed template can be mapped by properly restricting the action space $A^{2}(s)$.
where $v_{i_{j}}$ is a product node matched with the $j$-th node of the product pattern, and $v_{kl}$ is the $l$-th node of the $k$-th reactant pattern from $R$. To embed the reactant pattern nodes, we introduce a GNN $\mathrm{gnn}_{2}$ with the same architecture as $\mathrm{gnn}_{1}$. The action space $A^{3}(s)$ contains all possible atom mappings. We call an atom mapping between two nodes possible when the atom symbols of the nodes are the same and neither of the nodes has been previously mapped. The third phase ends when every node from the product pattern is mapped, resulting in a template that can be applied to the reaction center chosen in the first phase. The obtained reaction forms the terminal state $t$.
3.3 Training
We trained our RetroGFN with a modified version of Trajectory Balance Objective from [malkin2022trajectory], which for a trajectory $\tau=(s_{1},a_{1},s_{2},a_{2},...,s_{k},a_{k},t)$ is given with the formula:
The main difference from the original formulation comes from the fact that our RetroGFN is conditioned on the product from the initial state $s_{1}$. Therefore, for every initial state, we estimate the incoming flow separately using $F(s_{1})$ function which is essentially an index embedding $F(s)=E_{P}(s)\in\mathbb{R}$ that looks up the set of training products $P$ and returns a learnable scalar (note that we only evaluate $F(s)$ during training). As a backward policy $P_{B}(a|s)$, we use a uniform distribution over the possible actions that could lead to state $s$. The reward is an exponential reward of the form $R(x)=\mathrm{exp}(\beta f(x))$ where $f$ is a proxy for the desired property.
Reaction Feasiblity Guidance
. In RetroGFN, we use a reaction feasibility model as a proxy $f$. This way, we allow for exploration outside the limited reaction dataset as RetroGFN can assess the reaction feasibility via proxy $f$. The feasibility proxy can be a machine learning model that predicts the feasibility of a reaction or an indicator of whether the forward reaction prediction model was able to backtranslate the reaction $x$. In the main part of the paper, we evaluate the latter, while experiments with the former can be found in the Appendix B. Importantly, the forward model used during training was trained only on the training set of USPTO-50k, making it distinct from the model used in the round-trip evaluation.
3.4 Inference
$p(t)=\sum_{\tau:t\in\tau}\prod_{(s,a)\in\tau}P_{F}(a|s).$ (1)
To increase the accuracy of the estimation, we sample $K\cdot N$ trajectories. We leave the exploration of other estimation methods for future work. The details on the architecture and hyperparameters of both training and inference can be found in the A.
Table 1: Top-k round-trip accuracy results on USPTO-50k along with the standard deviation. The models are grouped into template-free, template-based, and semi-template models, respectively. The best results in every column are bolded. We observe that for $k>1$ our RetroGFN consistently outperforms other methods.
model top-1 top-3 top-5 top-10 top-20 top-50
MEGAN 88.2 81.4 77.0 68.8 58.3 41.8
RootAligned 94.5 85.3 79.0 67.8 53.3 27.1
Chemformer 92.6 58.9 42.0 24.2 12.9 5.3
GLN 93.4 87.6 84.2 77.8 67.5 47.5
MHNreact 91.9 85.6 81.0 72.1 60.1 38.5
LocalRetro 94.2 88.3 85.1 79.6 71.2 54.3
RetroKNN 94.0 87.8 84.3 78.7 70.0 53.1
Graph2Edits 92.6 82.8 75.9 61.5 42.8 21.3
GraphRetro 93.1 70.0 60.4 45.5 30.2 14.6
RetroGFN 92.8 88.8 86.1 81.4 74.9 63.6
Table 2: Top-k accuracy and mean reciprocal rank (mrr) results on USPTO-50k. The models are grouped into template-free, template-based, and semi-template models, respectively. The numbers in columns denote $k$ values. The best results in every column are bolded. We observe that for $k>3$ our RetroGFN achieves competitive results.
method mrr top-1 top-3 top-5 top-10 top-20 top-50
MEGAN 0.6231 48.7 72.4 79.5 86.8 91.0 93.5
RootAligned 0.6886 56.0 79.1 86.1 91.0 93.3 94.2
Chemformer 0.6312 55.0 70.9 73.7 75.4 75.9 76.0
GLN 0.6509 52.4 74.6 81.2 88.0 91.8 93.1
MHNreact 0.6350 50.8 72.7 79.6 86.3 90.0 92.3
LocalRetro 0.6587 52.0 76.6 84.6 91.1 94.9 96.7
RetroKNN 0.6837 55.7 77.9 85.5 91.7 94.7 96.5
Graph2Edits 0.6684 54.6 76.6 82.8 88.7 91.1 91.7
GraphRetro 0.6087 53.7 67.0 70.6 73.4 74.0 74.2
RetroGFN 0.6308 49.2 73.3 81.1 88.0 92.2 95.3
Table 3: Top-k round-trip accuracy on USPTO-MIT along with the standard deviation. The models are grouped into template-free, template-based, and semi-template models, respectively. The best results in every column are bolded. We observe that for $k>3$ our RetroGFN consistently outperforms other methods.
model top-1 top-3 top-5 top-10 top-20 top-50
MEGAN 83.2 75.8 71.3 63.5 54.1 39.4
RootAligned 85.9 76.6 70.6 60.7 48.1 24.8
Chemformer 82.8 53.5 38.9 22.8 12.2 5.0
GLN 84.2 77.5 73.6 66.9 57.1 39.4
MHNreact 83.3 76.5 72.2 64.0 52.6 32.7
LocalRetro 85.3 79.2 76.0 70.4 62.8 48.0
RetroKNN 85.3 77.5 72.9 62.5 42.9 18.6
Graph2Edits 83.9 74.3 67.9 55.5 39.4 20.3
GraphRetro 84.3 63.0 53.6 39.4 25.6 12.3
RetroGFN 83.9 78.8 76.0 71.4 65.8 56.2
Table 4: Top-k accuracy and mean reciprocal rank (mrr) results on USPTO-MIT. The models are grouped into template-free, template-based, and semi-template models, respectively. The best results in every column are bolded. We observe that our RetroGFN achieves competitive results.
method mrr top-1 top-3 top-5 top-10 top-20 top-50
MEGAN 0.4647 37.0 53.4 58.8 63.7 66.8 68.9
RootAligned 0.4960 40.2 56.8 61.7 66.1 68.5 69.5
Chemformer 0.4521 39.5 50.5 52.6 53.8 54.1 54.2
GLN 0.4625 37.2 52.9 57.7 62.7 65.2 66.3
MHNreact 0.4610 37.2 52.7 57.6 62.1 64.7 66.2
LocalRetro 0.4720 36.7 55.1 60.8 66.0 68.7 70.5
RetroKNN 0.4572 35.6 53.3 59.1 64.3 66.2 66.4
Graph2Edits 0.4795 38.9 55.2 60.1 64.0 65.6 66.1
GraphRetro 0.4461 39.1 49.2 51.8 53.7 54.3 54.6
RetroGFN 0.4590 35.4 53.5 59.3 64.7 67.8 70.0
4 Experiments
4.1 Setup
Datasets. We compared the considered methods on two datasets: USPTO-50k, a default choice for benchmarking retrosynthesis models, and USPTO-MIT, which we use as a generalization benchmark for models trained on USPTO-50k. We used commonly used splits for both datasets [coley2017retrosim, jin2017predicting]. We refined the USPTO-MIT to ensure there is no overlap between it and the USPTO-50k train split.
Retrosynthesis Models. We compared our RetroGFN to well-known and recent state-of-the-art retrosynthesis models; template-free: MEGAN [sacha2021molecule], RootAligned [zhong2022root] and Chemformer [irwin2022chemformer]; template-based: GLN [dai2019retrosynthesis], , MHNreact [seidl2021modern], LocalRetro [chen2021deep], RetroKNN [xie2023retroknn]; semi-template-based: Graph2Edits [graph2edit], GraphRetro [graphretro]. We used the wrappers of the original implementations and checkpoints from the Syntheseus repository [^1] https://github.com/microsoft/syntheseus. We used the evaluation procedure from Syntheseus that queries the model for 100 reactions, removes the duplicates, and truncates the list of reactions for every product to be no larger than 50. The same output was used both to calculate standard and round-trip metrics. All models were trained using the same train/validation/test splits.
Forward Model. A forward (reaction prediction) model takes a set of reactants as an input and outputs a set of possible products. As a backbone, we used a pre-trained Chemformer model from [irwin2022chemformer]. We fine-tuned two forward models: Chemformer-Eval, which was used to estimate the reaction feasibility in the round-trip accuracy (see Section 5), and Chemformer-Train, which guided RetroGFN during the training (see Section 3.3). Chemformer-Train was fine-tuned on the train split of USPTO-50k, while Chemformer-Eval used both the train and test split of USPTO-50k.
4.2 Results on USPTO-50k
The top-k round-trip accuracy results for the USPTO-50k dataset can be found in Table 1. Note that the forward model used during the training of RetroGFN was trained on a different data split than the one used for evaluation. We observe that for $k>1$ RetroGFN consistently outperforms all the models. The absolute and relative advantage of RetroGFN over the second-best model on top-k round-trip increases with $k$, which indicates two things. First, the model can return many diverse and feasible reactions while other models struggle to do that for $k>10$ (e.g., Chemformer). Second, while RetroGFN returns a lot of high-quality candidates, it does not rank them optimally. Therefore, its results on metrics that are sensitive to the quality of ranking, namely metrics with low $k$, are impoverished. We believe that improving the RetroGFN’s inference described in Section 3.4 may be a remedy for its performance with low $k$, but we leave it for future work.
In Table 2, we can find standard top-k accuracy results. Our method performs competitively with state-of-the-art single-step retrosynthesis models, especially for larger values of $k$, which are arguably more important for retrosynthesis search than $k=1$.
The good results of RetroGFN on standard metrics and its performance on round-trip accuracy evidence that one can improve the results on round-trip accuracy without sacrificing the performance on standard metrics, especially for larger values of k. Interestingly, the Pearson correlation between top-k accuracy and top-k round-trip for k=1 seems relatively high (corr=0.64, p-value=0.08) and increases with k(corr=0.85 and p-value=0.006 for k=50), indicating that there may be a space for jointly maximizing both metrics.
4.3 Generalization Results on USPTO-MIT
4.4 Diversity
We argue that the diversity of the reactants proposed by the model is captured by round-trip accuracy (Table 1 and 3), which counts the number of unique reactions that were predicted as feasible by some auxiliary model. Such a uniqueness-based notion of diversity is used in the GFlowNets literature [bengio2021flow], and we believe it is aligned with the retrosynthesis problem, where the reactions proposed for a single product are inherently similar (they share a significant number of atoms). However, to further assess the diversity of the models, we additionally reported the number of unique molecular scaffolds discovered among feasible reactions. Figure 4 shows that RetroGFN returns diverse sets of reactions, especially for larger k values.
Figure 4: A plot showing the diversity of the reactions proposed by various single-step retrosynthesis models. It shows the mean number of distinct molecular scaffolds observed in the top-k returned reactions that were predicted as feasible by a forward reaction prediction model (fine-tuned on USPTO-50k and USPTO-MIT, respectively). We observe that RetroGFN is able to return visibly more diverse reactions than other models.
4.5 Inference Time
As inference time is an important factor for retrosynthesis prediction, we report it in Table 5 for all the considered models for USPTO-50k and USPTO-MIT test sets. We observe a large variety of inference times. For USPTO-50k, it ranges from $0.39s$ per molecule for LocalRetro to $15.11s$ for Graph2Edits. Our RetroGFN has a moderate inference time: it ranks $7/10$ on USPTO-50k and $5/10$ on USPTO-MIT. The inference time can be reduced by optimizing the inference formulas from Section 3.4, which we leave for future work. We obtain almost 2x faster inference on USPTO-MIT because we set different inference parameters for this dataset (details in A). All times are reported using the Syntheseus repository [syntheseus] and V100 Nvidia GPUs.
Table 5: Table with average inference time per molecule in seconds for USPTO-50k and USPTO-MIT test sets.
method USPTO-50k USPTO-MIT
MEGAN 0.74 0.76
RootAligned 1.55 4.49
Chemformer 6.35 5.79
GLN 0.88 0.71
MHNreact 1.31 8.67
LocalRetro 0.39 0.43
RetroKNN 0.24 0.25
Grap2Edits 15.11 12.88
GraphRetro 8.52 4.3
RetroGFN 3.7 1.99
4.6 Multi-step Retrosynthesis
We evaluated single-step synthesis models in the multi-step search scenario using the Syntheseus repository [syntheseus]. We run each model on Monte Carlo Tree Search (MCTS) and Retro* [chen20retrostar] multi-step algorithms on 190 hard molecules from [chen20retrostar]. Each model is allowed to be called at most 600 times per target molecule and has the search time limited to $30$ minutes. Figure 5 shows the average number of calls required to solve a molecule and the number of non-overlapping ground-truth synthesis routes. Our RetroGFN achieves moderate results, being outperformed by models with higher top-k accuracy, e.g., LocalRetro, RootAligned or RetroKNN. According to [hartog2024investigations], the diversity of the single-step candidates and the quality of their ranking are important factors of multi-step search efficiency. The results from Table 1, 2 and Figure 4 suggest that RetroGFN outputs diverse candidates (great results for $k>10$), but struggles to rank them optimally (worse results for $k<3$) which diminishes its performance in the multi-step setting. We believe that improving the RetroGFN’s inference (Section 3.4) is an important future direction to enhance its real-world performance.
Figure 5: Multi-step search results on the Retro* Hard target set with different single-step models. Left: The number of calls until the first solution was found (or $\emptyset$ if a molecule was not solved). The orange line represents the median, the box represents the 25th and 75th percentile, the whiskers represent the 5th and 95th percentile, and points outside this range are shown as dots. Right: Approximate number of non-overlapping routes present in the search graph (tracked over the number of single-step model calls). The solid line represents the median, shaded area shows the 40th and 60th percentiles. On the right-hand side, we note the average time of solving the molecule.
4.7 Leveraging the Forward Model
In C, we study a simple model-agnostic way of leveraging the Chemformer-Train to maximize the results of the round-trip accuracy metric. While this approach significantly improves the round-trip accuracy results, it drastically decreases the standard top-k accuracy, especially for larger values of $k$. We leave the development of other methods of incorporating the Chemformer-Train model into the training pipeline for future work.